# GGHead: Fast and Generalizable 3D Gaussian Heads

### v1

@@ Abstract.
Learning 3D head priors from large 2D image collections is an important step towards high-quality 3D-aware human modeling. A core requirement is an efficient architecture that scales well to large-scale datasets and large image resolutions. Unfortunately, existing 3D GANs struggle to scale to generate samples at high resolutions due to their relatively slow train and render speeds, and typically have to rely on 2D superresolution networks at the expense of global 3D consistency. To address these challenges, we propose Generative Gaussian Heads (GGHead), which adopts the recent 3D Gaussian Splatting representation within a 3D GAN framework. To generate a 3D representation, we employ a powerful 2D CNN generator to predict Gaussian attributes in the UV space of a template head mesh. This way, GGHead exploits the regularity of the template’s UV layout, substantially facilitating the challenging task of predicting an unstructured set of 3D Gaussians. We further improve the geometric fidelity of the generated 3D representations with a novel total variation loss on rendered UV coordinates. Intuitively, this regularization encourages that neighboring rendered pixels should stem from neighboring Gaussians in the template’s UV space. Taken together, our pipeline can efficiently generate 3D heads trained only from single-view 2D image observations. Our proposed framework matches the quality of existing 3D head GANs on FFHQ while being both substantially faster and fully 3D consistent. As a result, we demonstrate real-time generation and rendering of high-quality 3D-consistent heads at $1024^{2}$ resolution for the first time.
@@
footnotetext: Project Website: https://tobias-kirschstein.github.io/gghead/
@@ 1. Introduction
A high-quality 3D generative model that can sample human heads with diverse geometry and appearance has to satisfy two important constraints: A high rendering resolution and strict 3D consistency of generated imagery. Once met, these properties enable exciting use cases such as single-image-to-3D reconstruction, 3D content creation, or 3D-consistent editing of images.
@@
A well-established approach are 3D Generative Adversarial Networks (3D GANs) that can learn unconditional 3D generation models from 2D images with corresponding camera poses. These methods have shown remarkable 3D-aware image synthesis capabilities, rapidly closing the quality gap to methods that directly generate 2D face images without view control. Under the hood, 3D GANs combine a differentiable rendering pipeline with a view-conditioned discriminator architecture to supervise 3D generation solely with 2D images. To obtain high-quality 3D samples with this approach, a high rendering resolution is crucial. A commonly chosen and successful differentiable rendering pipeline is Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021). However, so far, scaling the rendering resolution of NeRFs has been bottle-necked by the expensive ray marching procedure necessary to synthesize the final image. Several works have attempted to alleviate this limitation: EG3D (Chan et al., 2022) renders images at lower resolutions and then upsamples them using 2D superresolution, sacrificing 3D consistency in the process. GRAM (Deng et al., 2022) and GRAM-HD (Xiang et al., 2023) speed up ray marching by simplifying the 3D representation to a set of 2D manifolds, allowing cheap ray intersection tests. However, the simplistic 3D representation worsens rendering quality especially for side-views. Mimic3D (Chen et al., 2023) generates a pseudo-GT dataset from EG3D to supervise the 3D representation directly with patch-based rendering. Due to the inherent view-inconsistencies in the generated multi-view images from EG3D, only a perceptual loss can be used for supervision to gloss over the misalignment. Most recently, Trevithick et al. (2024) investigate how ray marching can be sped up with proposal networks, allowing training at $512^{2}$ resolution. Nevertheless, rendering speeds during both training and inference remain a significant challenge. For instance, currently, none of the existing methods are capable of natively generating and rendering images at full $512^{2}$ resolution in real-time. Since the creation of a 3D GAN involves several million render passes, this imposes a considerable computational burden on training.
@@ 2. Related Work
To obtain higher resolutions, some methods such as EG3D (Chan et al., 2022) apply a super-resolution network to the rendered features. While efficient, this approach can introduce unwanted artefacts in the form of 3D inconsistencies as well as low-resolution geometry. To alleviate this issue, several works focus on rendering in the pixel-space while maintaining fidelity and efficiency. Among them, Skorokhodov et al. (2022) propose to render patches instead of full images to reduce the computational cost. While this allows supervising a generative 3D representation without super-resolution, it impacts global consistency due to the discriminator’s limited field of view. Mimic3D (Chen et al., 2023) proposes an imitation strategy that forces the 3D representations to mimic the result of the 2D super-resolution branch. Similarly, GRAM-HD (Xiang et al., 2023) utilizes radiance manifolds and performs super-resolution on 2D manifolds instead of the rendered images. More recently, Trevithick et al. (2024) introduce a learning-based sampling strategy to effectively render every pixel and to model high-resolution geometry. Orthogonally to these work, our method aims to improve rendering speed by adopting 3D Gaussian Splatting as a 3D representation instead of NeRFs.
@@ 3.4.2. Gaussian Scale Regularization
(7) $\mathcal{L}_{reg}^{scale}=\|M_{scale}-\gamma_{scale}\|_{2}$
where $\gamma_{scale}=e^{-5}$ is a target scale specified to be just large enough that Gaussians would cover the template mesh.
@@ 3.5. Training Pipeline
where $\mathcal{L}_{adv}$ is the standard non-saturating GAN loss (Goodfellow et al., 2014) applied on the rendered 3D Gaussian representation. The discriminator $D$ is trained with R1 gradient regularization (Mescheder et al., 2018) with weight $1$. For the regularization weights, we choose $\lambda_{p}=0.1$ and $\lambda_{s}=0.05$. The opacity and UV total variation loss are initially turned off for training at $256^{2}$ resolution and are set to $\lambda_{o}=1$ and $\lambda_{uv}=100$ as soon as progressive growing is applied. We use a batch size of $32$ and learning rates of $0.02$ for both generator and discriminator using the Adam optimizer (Kingma and Ba, 2014). In total, we train our model for 25M images following Chan et al. (2022) which takes roughly 12 days on 2 RTX A6000 GPUs with 48G of VRAM each.
Figure 3. Generated Geometry: We show samples from our model with corresponding depth and normal maps. GGHead can generate fine geometric details such as eyeglasses and hair strands.
@@
4. Experimental Results
4.1. Dataset
We train our models on data from the FFHQ dataset (Karras et al., 2019b), which consists of 70k mostly frontal images of human faces. We use the preprocessed version of Chan et al. (2022) who aligned the images via facial landmarks, cropped them into $512^{2}$ resolution, and computed camera poses by resorting to a 3DMM fitting (Deng et al., 2019). We further follow EG3D’s processing pipeline to obtain a $1024^{2}$ resolution dataset of cropped and aligned face images with corresponding camera poses from the original FFHQ in-the-wild dataset. Finally, we employ the off-the-shelf background matting network MODNet (Ke et al., 2022) to replace the background with white pixels. That way, our models can focus their generation power solely on the actual 3D head. We conduct all our comparisons on this masked 2D face image dataset derived from FFHQ.
@@ 4.2. Baselines
For EG3D and Mimic3D, we use the official implementation and retrain their methods on our masked FFHQ dataset. Since Gaussian Shell Maps was originally proposed for animatable full body synthesis, we adapt their implementation and retrain it on our masked FFHQ dataset. For GRAM-HD, we use official model checkpoints to conduct qualitative and timing comparisons.
@@ 4.3. Qualitative Comparison
Figure 3 shows samples from our model alongside extracted depth and normal maps. The generated 3D representations exhibit a great amount of geometric detail and diversity. We further compare samples from GGHead to those of EG3D, Mimic3D, GRAM-HD and GSM in fig. 4. Our method produces renderings of great quality matching the current state-of-the-art at much higher speed. Interestingly, GSM synthesizes images with noticeable artifacts, indicating that generating 3D Gaussians alone is not sufficient for high-quality renderings. We attribute this to the design of GSM which disallows Gaussians from moving, severely restricting the expressiveness of the generated 3D representation.
@@
Table 1. Quantitative comparison: We compare our method to other 3D GANs on FFHQ at $512^{2}$ resolution. Our method fares favorably against other fully 3D-consistent methods and comes close to the performance of EG3D despite not using any 2D super-resolution. GGHead also produces more consistent 3D surfaces while simultaneously being considerably faster.
Method FID $\downarrow$ $\mathrm{PSNR}_{mv}\uparrow$ $\mathrm{SSIM}_{mv}\uparrow$
FFHQ@512 w/ SR EG3D 3.28 28.79 0.886
w/o SR GSM 28.19 25.53 0.895
Mimic3D 4.27 28.07 0.921
Ours 4.06 33.55 0.959
@@
4.4. Quantitative Comparison
We conduct quantitative comparisons with EG3D, Mimic3D, and Gaussian Shell Maps on the masked FFHQ dataset at $512^{2}$ resolution. Since there is no direct ground truth to evaluate the quality of generated 3D representations, we resort to measuring the Fréchet Inception Distance (FID) (Heusel et al., 2017). To do that, we generate 50k samples for each trained model using the training camera distribution and compare its statistics to the full dataset. The results are shown in table 1. Our method matches the performance of current state-of-the-art 3D GANs, obtaining slightly better FID scores than Mimic3D. It is only surpassed by EG3D, which employs a super-resolution module to achieve its rendering quality. To quantify the adverse effects of screen-space super-resolution on 3D consistency, we follow the procedure laid out by GRAM-HD (Xiang et al., 2023). For each method, we generate 30 images from different viewpoints of the same person, and then train the surface reconstruction method NeuS2 (Wang et al., 2023) on the generated images. We then measure the reconstruction error with PSNR and SSIM scores. The reasoning behind this is that view-inconsistencies in the images given to NeuS2 will lead to a higher reconstruction error. In practice, we observe that this metric also measures “surfaceness” of a 3D representation, since NeuS2 favors smooth and well-defined surfaces by design. We report the multi-view surface reconstruction results averaged over 30 randomly generated heads in table 1. GGHead leads to significantly better surface reconstructions than EG3D, underscoring its great 3D consistency.
4.5. Timing Comparison
@@ 4.6. Ablations
$R_{uv}$ Rendered Image $R_{uv}$ Rendered Image
without $\mathcal{L}_{uv}$ Ours
@@ 4.6.2. Effect of UV Total Variation Loss
Being fundamentally a 3D point representation, the 3D Gaussians lack the notion of a proper surface. Consequently, it can happen during training that Gaussians in the face area become too small, allowing other Gaussians from further back to shine through. This can create high frequency detail that satisfies the discriminator loss and looks plausible in static renderings. However, such a 3D representation remains flawed and is easily exposed in video renderings. Since we cannot supervise the time dimension during training, we instead employ our novel UV total variation loss. The effect can be seen in fig. 8. The regularization term effectively reduces the amount of Gaussians that shine through, considerably improving the surface quality of the generated 3D representations. As a side effect, the loss also resolves another failure case where lines of Gaussians would appear that mimic the appearance of hair strands but occur out of place. In table 3, we also study the effect of the UV total variation loss quantitatively and find that it slightly improves FID.
Table 3. Ablation study: We analyze the effect of different template meshes, different numbers of Gaussians, and our novel UV total variation loss.
[…]
256 Sphere Template 4.59
Plane Template 4.27
FLAME Template 4.74
Adjusted FLAME Template 4.03
512 65k Gaussians 6.39
262k Gaussians 4.28
+ $\mathcal{L}_{uv}$ (Ours) 4.06

### v2

@@ Abstract.
Learning 3D head priors from large 2D image collections is an important step towards high-quality 3D-aware human modeling. A core requirement is an efficient architecture that scales well to large-scale datasets and large image resolutions. Unfortunately, existing 3D GANs struggle to scale to generating samples at high resolutions due to their relatively slow train and render speeds, and typically have to rely on 2D superresolution networks at the expense of global 3D consistency. To address these challenges, we propose Generative Gaussian Heads (GGHead), which adopts the recent 3D Gaussian Splatting representation within a 3D GAN framework. To generate a 3D representation, we employ a powerful 2D CNN generator to predict Gaussian attributes in the UV space of a template head mesh. This way, GGHead exploits the regularity of the template’s UV layout, substantially facilitating the challenging task of predicting an unstructured set of 3D Gaussians. We further improve the geometric fidelity of the generated 3D representations with a novel total variation loss on rendered UV coordinates. Intuitively, this regularization encourages that neighboring rendered pixels should stem from neighboring Gaussians in the template’s UV space. Taken together, our pipeline can efficiently generate 3D heads trained only from single-view 2D image observations. Our proposed framework matches the quality of existing 3D head GANs on FFHQ while being both substantially faster and fully 3D consistent. As a result, we demonstrate real-time generation and rendering of high-quality 3D-consistent heads at $1024^{2}$ resolution for the first time.
Project Website: https://tobias-kirschstein.github.io/gghead
3D GAN, 3D head prior, 3D Gaussian Splatting
submissionid: 1054journal: TOGjournalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687686isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Reconstructionccs: Computing methodologies 3D imagingccs: Computing methodologies Adversarial learning
@@ 1. Introduction
A high-quality 3D generative model that can sample human heads with diverse geometry and appearance has to satisfy two important constraints: A high rendering resolution and strict 3D consistency of generated imagery. Once met, these properties enable exciting use-cases such as single-image-to-3D reconstruction, 3D content creation, or 3D-consistent editing of images.
@@
A well-established approach are 3D Generative Adversarial Networks (3D-GANs) that can learn unconditional 3D generation models from 2D images with corresponding camera poses. These methods have shown remarkable 3D-aware image synthesis capabilities, rapidly closing the quality gap to methods that directly generate 2D face images without view control. Under the hood, 3D GANs combine a differentiable rendering pipeline with a view-conditioned discriminator architecture to supervise 3D generation solely with 2D images. To obtain high-quality 3D samples with this approach, a high rendering resolution is crucial. A commonly chosen and successful differentiable rendering pipeline is Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021). However, so far, scaling the rendering resolution of NeRFs has been bottle-necked by the expensive ray marching procedure necessary to synthesize the final image. Several works have attempted to alleviate this limitation: EG3D (Chan et al., 2022) renders images at lower resolutions and then upsamples them using 2D superresolution, sacrificing 3D consistency in the process. GRAM (Deng et al., 2022) and GRAM-HD (Xiang et al., 2023) speed up ray marching by simplifying the 3D representation to a set of 2D manifolds, allowing cheap ray intersection tests. However, the simplistic 3D representation worsens rendering quality especially for side-views. Mimic3D (Chen et al., 2023) generates a pseudo-GT dataset from EG3D to supervise the 3D representation directly with patch-based rendering. Due to the inherent view-inconsistencies in the generated multi-view images from EG3D, only a perceptual loss can be used for supervision to gloss over the misalignment. Most recently, Trevithick et al. (2024a) investigate how ray marching can be sped up with proposal networks, allowing training at $512^{2}$ resolution. Nevertheless, rendering speeds during both training and inference remain a significant challenge. For instance, currently, none of the existing methods are capable of natively generating and rendering images at full $512^{2}$ resolution in real-time. Since the creation of a 3D GAN involves several million render passes, this imposes a considerable computational burden on training.
@@ 2. Related Work
To obtain higher resolutions, some methods such as EG3D (Chan et al., 2022) apply a super-resolution network to the rendered features. While efficient, this approach can introduce unwanted artefacts in the form of 3D inconsistencies as well as low-resolution geometry. To alleviate this issue, several works focus on rendering in the pixel-space while maintaining fidelity and efficiency. Among them, Skorokhodov et al. (2022) propose to render patches instead of full images to reduce the computational cost. While this allows supervising a generative 3D representation without super-resolution, it impacts global consistency due to the discriminator’s limited field of view. Mimic3D (Chen et al., 2023) proposes an imitation strategy that forces the 3D representations to mimic the result of the 2D super-resolution branch. Similarly, GRAM-HD (Xiang et al., 2023) utilizes radiance manifolds and performs super-resolution on 2D manifolds instead of the rendered images. More recently, Trevithick et al. (2024b) introduce a learning-based sampling strategy to effectively render every pixel and to model high-resolution geometry. Orthogonally to these work, our method aims to improve rendering speed by adopting 3D Gaussian Splatting as a 3D representation instead of NeRFs.
@@ 3.3.1. Gaussian Positions
3.3.2. Gaussian Scales
In similar spirit, we limit the range of predicted Gaussian scales:
(6) $\mathcal{G}_{scale}^{i}$ $\leftarrow\textsc{exp}\left(-s_{max}-\textsc{softplus}(-(G_{scale}^{i}-s_{init})-s_{max})\right)$
which ensures that scales are within $[0,e^{-s_{max}}]$ and initially default to $e^{-s_{init}}$ when the network predicts zeros. We set $s_{max}=3$ and $s_{init}=5$ which initializes Gaussian scales with $\sim$7mm roughly covering the template mesh, and limits them from exceeding $\sim$5cm.
@@ 3.4.2. Gaussian Scale Regularization
(8) $\mathcal{L}_{reg}^{scale}=\|M_{scale}\|_{2}$
In combination with eq. 6 this regularization pushes Gaussian scales to be close to $\sim$7mm which is just large enough that Gaussians would cover the template mesh.
@@ 3.4.3. Gaussian Opacity Regularization
$R_{uv}$ Rendered Image $R_{uv}$ Rendered Image
without $\mathcal{L}_{uv}$ Ours
@@ 3.4.4. UV Total Variation Loss
AFHQ-M FFHQ-M
Figure 4. Generated Geometry: We show samples from our models trained on FFHQ-M and AFHQ-M on $512^{2}$ resolution with corresponding depth and normal maps. GGHead can generate fine geometric details such as eyeglasses and hair strands.
@@ 3.5. Training Pipeline
where $\mathcal{L}_{adv}$ is the standard non-saturating GAN loss (Goodfellow et al., 2014) applied on the rendered 3D Gaussian representation. The discriminator $D$ is trained with R1 gradient regularization (Mescheder et al., 2018) with weight $1$. For the regularization weights, we choose $\lambda_{p}=0.1$ and $\lambda_{s}=0.05$. The opacity and UV total variation loss are initially turned off for training at $256^{2}$ resolution and are set to $\lambda_{o}=1$ and $\lambda_{uv}=100$ as soon as progressive growing is applied. We use a batch size of $32$ with a learning rate of $0.0025$ for the generator and $0.002$ for the discriminator discriminator using the Adam optimizer (Kingma and Ba, 2014). We also incorporate the speed improvements by Durvasula et al. (2023) for faster differentiable rasterization of Gaussians. In total, we train our model for 25M images following Chan et al. (2022) which takes roughly 12 days on 2 RTX A6000 GPUs with 48G of VRAM each.
4. Experimental Results
@@
4.1. Datasets
We train our models on data from the FFHQ dataset (Karras et al., 2019b), which consists of 70k mostly frontal images of human faces. We use the preprocessed version of Chan et al. (2022) who aligned the images via facial landmarks, cropped them into $512^{2}$ resolution, and computed camera poses by resorting to a 3DMM fitting (Deng et al., 2019). We further follow EG3D’s processing pipeline to obtain a $1024^{2}$ resolution dataset of cropped and aligned face images with corresponding camera poses from the original FFHQ in-the-wild dataset. Finally, we employ the off-the-shelf background matting network MODNet (Ke et al., 2022) to replace the background with white pixels yielding a masked version of FFHQ denoted as FFHQ-M. That way, our models can focus their generation power solely on the actual 3D head. We also apply the same masking strategy to the AFHQv2 Cats dataset (Choi et al., 2020) yielding AFHQ-M.
@@ 4.2. Baselines
For EG3D and Mimic3D, we use the official implementation and retrain their methods on the masked FFHQ and AFHQ datasets. Since Gaussian Shell Maps was originally proposed for animatable full body synthesis, we adapt their implementation and retrain it on the FFHQ-M dataset. For GRAM-HD, we use official model checkpoints to conduct qualitative and timing comparisons.
Table 1. Quantitative comparison: We compare our method to other 3D GANs on FFHQ and AFHQ at $512^{2}$ resolution. FFHQ-M and AFHQ-M denote the resepective datasets with the background removed. Our method fares favorably against other fully 3D-consistent methods and comes close to the performance of EG3D despite not using any 2D super-resolution. GGHead also produces more consistent 3D surfaces while simultaneously being considerably faster.
FFHQ-M FFHQ AFHQ-M
[…]
w/ SR EG3D 3.28 28.74 0.899 4.70 2.82
w/o SR GSM 28.19 26.40 0.919 - -
Mimic3D 4.27 30.95 0.949 5.37 4.48
Ours 4.06 33.29 0.964 5.15 3.40
4.3. Quantitative Comparison
We conduct quantitative comparisons with EG3D, Mimic3D, and Gaussian Shell Maps on the FFHQ and AFHQv2 Cats datasets at $512^{2}$ resolution. For comparisons on the unmasked FFHQ dataset, we extend our generator to predict 3 additional channels to produce a background image and simply blend that with the rendered Gaussian head. For AFHQ-M, we take a model trained on FFHQ-M and finetune it on the much smaller cats dataset following Chan et al. (2022). Since there is no direct ground truth to evaluate the quality of generated 3D representations, we resort to measuring the Fréchet Inception Distance (FID) (Heusel et al., 2017). To do that, we generate 50k samples for each trained model using the training camera distribution and compare its statistics to the full dataset. The results are shown in table 1. Our method matches the performance of current state-of-the-art 3D GANs, obtaining slightly better FID scores than Mimic3D. It is only surpassed by EG3D, which employs a super-resolution module to achieve its rendering quality. To quantify the adverse effects of screen-space super-resolution on 3D consistency, we follow the procedure laid out by GRAM-HD (Xiang et al., 2023). For each method, we generate 30 images from different viewpoints of the same person, and then train the surface reconstruction method NeuS2 (Wang et al., 2023) on the generated images. We then measure the reconstruction error with PSNR and SSIM scores. The reasoning behind this is that view-inconsistencies in the images given to NeuS2 will lead to a higher reconstruction error. In practice, we observe that this metric also measures “surfaceness” of a 3D representation, since NeuS2 favors smooth and well-defined surfaces by design. We report the multi-view surface reconstruction results averaged over 100 randomly generated heads in table 1. GGHead leads to significantly better surface reconstructions than EG3D, underscoring its great 3D consistency.
@@ 4.4. Qualitative Comparison
Figure 4 shows samples from our model alongside extracted depth and normal maps for both FFHQ-M and AFHQ-M. The generated 3D representations exhibit a great amount of geometric detail and diversity. We further compare samples from GGHead to those of EG3D, Mimic3D, GRAM-HD and GSM in fig. 5. Our method produces renderings of great quality matching the current state-of-the-art at much higher speed. Interestingly, GSM synthesizes images with noticeable artifacts, indicating that generating 3D Gaussians alone is not sufficient for high-quality renderings. We attribute this to the design of GSM which disallows Gaussians from moving, severely restricting the expressiveness of the generated 3D representation.
@@
4.5. Compute Resource Requirements
Finally, we record GPU memory requirements and the total training time in table 3. Our method is significantly more memory-efficient than other 3D GANs and finishes training noticeably faster.
@@ 4.5. Compute Resource Requirements
Table 3. Compute Resource Requirements: We measure the GPU memory consumption during training on a single RTX A6000 GPU (48GB VRAM). Out-of-memory is denoted as OOM. Our method is significantly more memory-efficient and therefore scales better with larger batch sizes. We also record the total training time in equivalent days on a single RTX A6000.
Batch Size
Method 4 8 16 32 Train time
Mimic3D 20.4 GB 38.7 GB OOM OOM 80.2 days
EG3D 13.4 GB 25.1 GB OOM OOM 39.0 days
Ours 8.6 GB 12.8 GB 19.4 GB 35.0 GB 23.4 days
Table 4. Ablation study: We analyze the effect of regularization, different template meshes, different numbers of Gaussians, and our novel UV total variation loss when training our model on FFHQ-M.
[…]
256 Regular-izations w/o $\mathcal{L}_{reg}^{scale}$ collapse
w/o $\mathcal{L}_{reg}^{pos}$ 10.92
w/o zero-init layer $\mathcal{Z}$ 6.71
Template Sphere Template 4.59
Plane Template 4.27
FLAME Template 4.74
Adjusted FLAME Template 4.03
512 Gaussians 65k Gaussians 6.39
262k Gaussians 4.28
+ $\mathcal{L}_{uv}$ (Ours) 4.06
@@ 4.6. Ablations
4.6.1. Effect of Regularization
In the upper part of table 4, we analyze the impact of the scale and position regularization terms $\mathcal{L}_{reg}^{scale}$ and $\mathcal{L}_{reg}^{pos}$, as well as the usefulness of the zero-initialized CNN layer $\mathcal{Z}$ that enforces small position offsets predictions in the beginning. We find that all measures improve training stability which translates to a better FID score. Without scale regularization, predicted Gaussians even disappear entirely leading to total training collapse.
@@ 4.6.3. Effect of UV Total Variation Loss
Being fundamentally a 3D point representation, the 3D Gaussians lack the notion of a proper surface. Consequently, it can happen during training that Gaussians in the face area become too small, allowing other Gaussians from further back to shine through. This can create high frequency detail that satisfies the discriminator loss and looks plausible in static renderings. However, such a 3D representation remains flawed and is easily exposed in video renderings. Since we cannot supervise the time dimension during training, however, we instead employ our novel UV total variation loss. The effect can be seen in fig. 3. The regularization term effectively reduces the amount of Gaussians that shine through, considerably improving the surface quality of the generated 3D representations. As a side effect, the loss also resolves another failure case where lines of Gaussians would appear that mimic the appearance of hair strands but occur out of place. In table 4, we also study the effect of the UV total variation loss quantitatively and find that it slightly improves FID.